A string of mass 3 grams has a lenght of 60cm.What must be the tension in the string so that when vibrating transversely it's first overtone has a frequency 200per sec


1
Expert's answer
2021-03-10T17:16:45-0500
"f=\frac{1}{2l}\sqrt{\frac{Tl}{m}}\\
200=\frac{1}{2(0.6)}\sqrt{\frac{0.6T}{0.003}}\\T=288\ N"